OpenStudy (anonymous):

HELPPP! Solve: 6(x-5)-3x=-9

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

6(x-5)-3x=-9 => 6x-30-3x=-9 => 3x-30(+30)=-9(+30) => 3x=21 => 3x/3=21/3 => x=7

OpenStudy (anonymous):

thank you!